# On the Mistaken Assumption of Interchangeable Deep Reinforcement Learning Implementations

### v1

@@ Abstract
Index Terms:
reinforcement learning, differential testing
**footnotetext: Corresponding author.footnotetext: Shenzhen Campus of Sun Yat-sen University.footnotetext: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ III-A Literature Review
In total, we identified ten papers from NeurIPS, six papers from ICML, and seven papers from other conferences under the assumption of interchangeable algorithm implementations. Examples of studies conforming to the first aforementioned scenario are the works by Beukman et al. [28], Chiappa et al. [29], and Gerstgrasser et al. [30], where alternate implementations were used over the original for the SAC [31], PPO, and DQN algorithms respectively. Examples of studies conforming to the second aforementioned scenario are the works by Raghunath et al. [17] and Wolk et al. [18]. Raghunath et al. [17] integrated their study with the SAC algorithm in a single-agent setting as well as a multi-agent setting to demonstrate effectiveness. However, Raghunath et al. used different SAC implementations for both settings, assuming that they were consistent and thus, interchangeable. Given that their experiments show marginal differences between both single-agent and multi-agent settings [17, Figure 2], if the SAC implementations were not interchangeable, a change in SAC implementation could easily flip the experimental outcomes. Similarly, Wolk et al. [18] integrated their study with the PPO algorithm in multiple settings. However, Wolk et al. used different PPO implementations for some settings, assuming they were interchangeable. This potentially invalidates their comparisons, especially for those with marginal differences [18, Table 2].
@@ IV-B Comparing DRL Performance
Firstly, we incorporated an environment-wise one-way ANOVA to compare the effect of PPO implementation on $MeanReward_{100}$. With a null hypothesis of equal implementation means, rejecting it (i.e., p-value $<$ 0.05) translates to a statistically significant difference in means. However, (1) a one-way ANOVA does not pinpoint outliers, (2) with large amounts of data, minor effects can easily cause statistically significant differences, and (3) statistical insignificance does not imply absence of effect [45, 25]. Thus, we decided to include additional techniques for broader perspectives.
@@
[…]
@@ V Results
    Reject null hypothesis of equal implementation means if p-value $<$ 0.05.
@@ V-A Prevalence of Implementation Discrepancies
[…]
@@ Acknowledgments
    Reject null hypothesis of equal implementation means if p-value $<$ 0.05.
@@
    Reject null hypothesis of equal implementation means if p-value $<$ 0.05.

### v2

@@ Abstract
**footnotetext: Corresponding author.footnotetext: Shenzhen Campus of Sun Yat-sen University.footnotetext: This work was published in the 2025 IEEE/ACM 47th International Conference on Software Engineering (ICSE) [1] and is available in IEEE Xplore.footnotetext: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ III-A Literature Review
In total, we identified ten papers from NeurIPS, six papers from ICML, and seven papers from other conferences under the assumption of interchangeable algorithm implementations. Examples of studies conforming to the first aforementioned scenario are the works by Beukman et al. [29], Chiappa et al. [30], and Gerstgrasser et al. [31], where alternate implementations were used over the original for the SAC [32], PPO, and DQN algorithms respectively. Examples of studies conforming to the second aforementioned scenario are the works by Raghunath et al. [18] and Wolk et al. [19]. Raghunath et al. [18] integrated their study with the SAC algorithm in a single-agent setting as well as a multi-agent setting to demonstrate effectiveness. However, Raghunath et al. used different SAC implementations for both settings, assuming that they were consistent and thus, interchangeable. Given that their experiments show marginal differences between both single-agent and multi-agent settings [18, Figure 2], if the SAC implementations were not interchangeable, a change in SAC implementation could easily flip the experimental outcomes. Similarly, Wolk et al. [19] integrated their study with the PPO algorithm in multiple settings. However, Wolk et al. used different PPO implementations for some settings, assuming they were interchangeable. This potentially invalidates their comparisons, especially for those with marginal differences [19, Table 2].
@@ IV-B Comparing DRL Performance
Firstly, we incorporated an environment-wise one-way ANOVA to compare the effect of PPO implementation on $MeanReward_{100}$. With a null hypothesis of equal implementation means, rejecting it (i.e., p-value << 0.05) translates to a statistically significant difference in means. However, (1) a one-way ANOVA does not pinpoint outliers, (2) with large amounts of data, minor effects can easily cause statistically significant differences, and (3) statistical insignificance does not imply absence of effect [46, 26]. Thus, we decided to include additional techniques for broader perspectives.
@@
Here, we directly compute the POI of implementation $X$ over implementation $Y$ with $N$ trials each, on environment $m$. The primary assumption is that for $X$ to be better than $Y$, $X$ has to outperform $Y$ sufficiently often. The final POI of $X$ over $Y$, $P(X>Y)$, is subsequently attained by calculating the POI for all environments $\frac{1}{M}\sum_{m=1}^{M}P(X_{m}>Y_{m})$. Post-SBCI, the point estimates and confidence bands attained from this metric are further analyzed with the Neyman-Pearson testing criterion [48, 49] for statistical significance and meaningfulness. Statistical significance where $P(X>Y)>0.5\;\wedge\;0.5\notin CI$ rules out the effect of noise on the point estimates. On the other hand, statistical meaningfulness where $CI_{upper}>0.75$ ensures that $X$ outperforms $Y$ often enough. X is considered to be better than Y if both aforementioned criteria are met. Thus, the POI metric directly measures the likelihood that an implementation $X$ outperforms implementation $Y$ on a random environment, complementing the performance profiles.
@@ V Results
    Reject null hypothesis of equal implementation means if p-value << 0.05.
@@ V-A Prevalence of Implementation Discrepancies
Figure 5 shows that the high-performing group is objectively better than the low-performing group. In particular, the pairwise POI between the two groups is both statistically significant, where $P(X>Y)>0.5\;\wedge\;0.5\notin CI$, and meaningful, where $CI_{upper}>0.75$. From the pairwise comparisons, it is also observed that Stable Baselines3, CleanRL, and Baselines are on par with each other, with the pairwise comparisons among themselves being neither statistically significant nor meaningful. However, as shown in Table II, when considering individual environments, we found statistically significant discrepancies. We later determined these discrepancies to be caused by implementation inconsistencies in RQ2.
@@ Acknowledgments
    Reject null hypothesis of equal implementation means if p-value << 0.05.
@@
    Reject null hypothesis of equal implementation means if p-value << 0.05.